# Regularity of Einstein $5$-manifolds via $4$-dimensional gap theorems

### v1

@@ 1.2.4. Singular set of noncollapsed limits of Einstein $n$-manifolds
Let $(M_{i},g_{i},p_{i})_{i\in\mathbb{N}}$ be a sequence of pointed $n$-manifolds with $|{\rm Ric}_{g_{i}}|\leqslant n-1$ and ${\rm Vol}(B_{1}(p_{i}))\geqslant v>0$, and let $(X,d,p)$ be a Gromov-Hausdorff limit. Fix $x\in X$. Assume that every tangent cone at $x$ splits off an $\mathbb{R}^{n-5}$-factor and that at least one tangent cone at $x$ splits off an $\mathbb{R}^{n-4}$-factor. Then the tangent cone at $x$ is unique isometric to $\mathbb{R}^{n-4}\times\mathbb{R}^{4}/\Gamma$ for some $\Gamma\subset O(4)$.
@@ 1.3.1. Dimension 5
Motivated by the Calabi–Yau singular models [HN, Sze19], it is reasonable to expect that singularities of the form $\mathbb{R}^{2k}\times\mathbb{R}^{4}/\Gamma$ in Einstein limits can be isolated; the same may plausibly hold for $\mathbb{R}\times\mathbb{R}^{4}/\Gamma$. The question is whether they can accumulate without forming curves. See also [Sze21] in the context of minimal surfaces.
@@
Motivated by the Calabi–Yau models [Li19, CR23, Sze19, Sze20a], with tangent cones of the form $\mathbb{R}^{2k}\times\mathbb{R}^{4}/\Gamma$, a natural question is the following one.
Question 1.26.
@@ 1.3.2. Higher dimensions
In higher dimensions, we have less evidence, but based on our $5$-dimensional results, we ask natural questions.
Question 1.27 (Higher-dimensional isolation of spherical orbifolds).
@@
Question 1.30 (Codimension-$5$ conjecture).
@@
Towards a proof in all dimensions. The present paper essentially proves the iteration from n1=4n-1=4 to $n=5$ of a plausible, long-term strategy in order to answer Question 1.30 affirmatively by an iteration on the dimension. Assume that noncollapsed limits of (n1)(n-1)-dimensional Einstein metrics have codimension $5$ regularity in the sense of Question 1.30.
@@ 1.4.2. Regularity of Einstein $5$-manifolds
Refined rectifiability. The isolation theorem implies that $1$-symmetry at a given scale propagates to uniform symmetry across all smaller scales. Using the cone-splitting principle and $\varepsilon$-regularity, this uniform symmetry allows the construction of almost splitting maps whose restrictions to the singular set are bi-Lipschitz onto intervals. A countable selection of such maps yields a covering of $\mathcal{S}^{1}$ by bi-Lipschitz images of subsets of $\mathbb{R}$, thereby establishing a refined rectifiability of the singular set without discarding any exceptional subsets. Moreover, we show that along each curve with constant volume density, the tangent cones remain identical. The proof is given in Section 4.3.
@@ 1.5. Acknowledgements
The first author is grateful to Tobias Colding for suggesting the uniqueness problem and for his patient guidance and many insightful discussions. The first author also thanks Wenshuai Jiang for his support and helpful conversations. During this project, Yiqi Huang was partially supported by a Simons Dissertation Fellowship and by NSF Grant DMS–2405393, and Tristan Ozuch was partially supported by the National Science Foundation under grant DMS–2405328.
@@ 4.1. Preliminaries
In [CJN21] and [JN21], the finiteness of the measure $\mathcal{H}^{n-4}(\mathcal{S}\cap B_{1}(p))\leqslant C(n,v)$ is proven as well as a structure theorem for the singular set $\mathcal{S}$:
@@
    If there exists a $\delta$-splitting function on $B_{2}(p)$, then $d_{GH}(B_{1}(p),B_{1}(0,x^{\prime}))\leqslant\varepsilon$ for $(0,x^{\prime})\in\mathds{R}\times X$.
@@ 4.3. Refined rectifiability
Let $X^{5}$ be a noncollapsed Ricci-bounded limit space. Recall that in [CJN21, JN21], it is proven that the top stratum $\mathcal{S}^{1}\setminus\mathcal{S}^{0}$ is contained in a countable union of bi-Lipschitz images from subsets of $\mathds{R}$, modulo some $\mathcal{H}^{1}$-measure zero set. In this section, building on the uniqueness Theorem, we will improve the bi-Lipschitz structure by showing that there is no need to discard any measure zero subset, thereby establishing Theorem 1.16. For related work on uniqueness and the singular set we refer to [CM16, NV17b, HJ23, HJ24, HJ25, FL25b].
@@
The key result is the following uniform symmetry propagation, coming from our gap Theorem 1.1.
@@
With this uniform propagation of symmetry, we construct a local bi-Lipschitz map around $\mathcal{S}^{1}_{\Gamma}$ using $\varepsilon$-splitting functions as in [JN21].
@@
For any $\varepsilon>0$ and any $x\in B_{1}(p)\cap\mathcal{S}^{1}_{\Gamma}$, there exists some function $u$ defined on $B_{r_{x}}(x)$ for some radius $r_{x}$ such that $u$ is $(1+\varepsilon)$-Lipschitz on $B_{r_{x}}(x)\cap\mathcal{S}^{1}_{\Gamma}$: for any $y,z\in B_{r_{x}}(x)\cap\mathcal{S}^{1}_{\Gamma}$
$\Big||u(y)-u(z)|-d(y,z)\Big|\leqslant\varepsilon d(y,z).$ (4.9)
@@
Let $\varepsilon>0$ be given and $\eta>0$. Since $x\in\mathcal{S}^{1}_{\Gamma}$, by rescaling we may assume that $d_{GH}(B_{2}(x),B_{2}(0,x^{\prime}))\leqslant 2\delta$ with $\delta\leqslant\delta_{0}(v,\varepsilon)$ in Theorem 4.10, where $(0,x^{\prime})$ is the vertex of the cone $\mathds{R}\times C(\mathbb{S}^{3}/\Gamma)$. By Theorem 4.10, there exists some $\varepsilon$-splitting function $u$ defined on $B_{1}(x)$. We will prove that this function $u$ satisfies the bi-Lipschitz condition.
For any $y\in B_{\delta}(x)\cap\mathcal{S}^{1}_{\Gamma}$, by triangle inequality, we have
$d_{GH}(B_{2}(y),B_{2}(0,x^{\prime}))\leqslant d_{GH}(B_{2}(y),B_{2}(x))+d_{GH}(B_{2}(x),B_{2}(0,x^{\prime}))\leqslant 3\delta.$
By Proposition 4.13, we have $d_{GH}(B_{r}(y),B_{r}(0,x^{\prime}))\leqslant\eta r$ for any $r\in(0,1]$, provided further $\delta\leqslant\delta(\eta)$. Let $\mathcal{L}_{y,r}$ be the image of $\mathds{R}\times\{x^{\prime}\}$ under the corresponding Gromov-Hausdorff map at $y$ at the scale $r$. Then by cone-splitting Lemma 4.7 and $\varepsilon$-regularity Theorem 4.8, for any $r\leqslant 1/4$, $\varepsilon>0$ and $\eta\leq\eta(\varepsilon)$ small, we have
$B_{r}(y)\cap\mathcal{S}^{1}_{\Gamma}\subset B_{\varepsilon r}(\mathcal{L}_{y,2r}).$ (4.10)
Claim: The function $u:B_{\varepsilon r}(\mathcal{L}_{y,2r})\to\mathds{R}$ is a $10\varepsilon r$-Gromov-Hausdorff map for any $r\leqslant 1/4$.
We can prove (4.9) by assuming the claim. In fact, pick any $y,z\in B_{\delta}(x)\cap\mathcal{S}^{1}_{\Gamma}$ with $d(y,z)=r/2\leqslant 1/8$. Since $u$ is a $10\varepsilon r$-Gromov-Hausdorff map on $B_{r}(y)\cap\mathcal{S}^{1}_{\Gamma}$, then
$\Big||u(y)-u(z)|-d(y,z)\Big|\leqslant 10\varepsilon r=20\varepsilon d(y,z).$
Then we finish the proof of (4.9) by taking $\varepsilon$ as small as needed.
It remains to prove the claim. Pick $y\in B_{\delta}(x)$ and $r\leqslant 1/4$. Since $d_{GH}(B_{r}(y),B_{r}(0,x^{\prime}))\leqslant\eta r$, then by Theorem 4.10, there exists a $\varepsilon$-splitting function $u_{y,r}$ defined on $B_{r}(y)$. Moreover, $u_{y,r}:\mathcal{L}_{y,2r}\cap B_{r}(y)\to\mathds{R}$ is a $2\varepsilon r$-Gromov-Hausdorff map. By translation we can assume $u(y)=u_{y,r}(y)$.
Note that if $\eta\leqslant\eta(\varepsilon)$ is chosen small, then up to a multiplicative constant $\pm 1$, we have $u_{y,r}$ and $u$ are $\varepsilon r$-close in $B_{2r}(y)$:
$\sup_{B_{2r}(y)}|\pm u_{y,r}-u|\leqslant\varepsilon r.$ (4.11)
To prove (4.11), we use a contradiction argument as follows. By the convergence of splitting function, Proposition 4.11, the contradiction sequences for $u$ and $u_{y,r}$, both converge to the coordinate function in $\mathds{R}\subset\mathds{R}\times C(\mathbb{S}^{3}/\Gamma)$ uniformly in compact subsets of $B_{2r}(y)$. This proves (4.11). Since $u_{y,r}$ is a $2\varepsilon r$-Gromov-Hausdorff map on $\mathcal{L}_{y,2r}\cap B_{r}(y)$, then it follows that $u$ is a $10\varepsilon r$-Gromov-Hausdorff map on $\mathcal{L}_{y,2r}\cap B_{r}(y)$. This completes the proof by taking $r_{x}=\delta$. ∎
Therefore, we can prove the refined rectifiability of the singular set $\mathcal{S}$.
@@
Since we have the uniform lower bound of ${\rm Vol}(B_{1}(x))$ for $x\in B_{1}(p)$, the index set $I$ is finite. For each $x\in\mathcal{S}^{1}_{\Gamma_{i}}\cap B_{1}(p)$, by Proposition 4.14, there exists some radius $r_{x}$ such that $B_{r_{x}}(x)\cap\mathcal{S}^{1}_{\Gamma_{i}}$ is bi-Lipschitz to a subset of $\mathds{R}$. Since $X$ is separable, there exist countably many subsets $Z_{j}\in\mathcal{S}^{1}$ such that each $Z_{j}$ is bi-Lipschitz to some subset of $\mathds{R}$ with $\mathcal{S}^{1}=\bigcup_{j\in J}Z_{j}$. The proof is completed by noting that $\mathcal{S}^{0}$ is also countable. ∎

### v2

@@ 1.2.4. Singular set of higher dimensional noncollapsed limits
Let $(M_{i},g_{i},p_{i})_{i\in\mathbb{N}}$ be a sequence of pointed $n$-manifolds with $|{\rm Ric}_{g_{i}}|\leqslant n-1$ and ${\rm Vol}(B_{1}(p_{i}))\geqslant v>0$, and let $(X,d,p)$ be a Gromov-Hausdorff limit. Fix $x\in X$. Assume that every tangent cone at $x$ splits off an $\mathbb{R}^{n-5}$-factor and that at least one tangent cone at $x$ splits off $\mathbb{R}^{n-4}$. Then the tangent cone at $x$ is unique isometric to $\mathbb{R}^{n-4}\times\mathbb{R}^{4}/\Gamma$ for some $\Gamma\subset O(4)$.
Corollary 1.24.
Let $(X,d)$ be a noncollapsed limit of $6$-manifolds with $|{\rm Ric}|\leqslant 5$. Let $p=\gamma(0)$ be a point lying on an open unit-speed geodesic $\gamma:(-\varepsilon,\varepsilon)\to X$ for some $\varepsilon>0$. Assume that one tangent cone at $p$ splits off $\mathbb{R}^{n-4}$. Then, there exists $\Gamma\subset O(4)$ such that for any $s\in(-\varepsilon,\varepsilon)$, the tangent cone at $\gamma(s)$ is unique and isometric to $\mathbb{R}^{2}\times\mathbb{R}^{4}/\Gamma$.
@@ 1.3.1. Dimension 5
Motivated by the Calabi-Yau singular models [HN, Sze19], it is reasonable to expect that singularities of the form $\mathbb{R}^{2k}\times\mathbb{R}^{4}/\Gamma$ in Einstein limits can be isolated; the same may plausibly hold for $\mathbb{R}\times\mathbb{R}^{4}/\Gamma$. The question is whether they can accumulate without forming curves. See also [Sze21] in the context of minimal surfaces.
@@
Motivated by the Calabi-Yau models [Li19, CR23, Sze19, Sze20a], with tangent cones of the form $\mathbb{R}^{2k}\times\mathbb{R}^{4}/\Gamma$, a natural question is the following one.
Question 1.29.
@@ 1.3.2. Higher dimensions
In higher dimensions, we have less evidence, but based on our $5$-dimensional results, we ask natural questions of isolation and gap theorems.
Question 1.30 (Higher-dimensional isolation of spherical orbifolds and gap theorem).
@@
Question 1.31.
For every $n\geq 4,v>0$, does there exist $\varepsilon_{0}=\varepsilon_{0}(n,v)>0$ such that if $(M_{o}^{n},g_{o})$ with ${\rm Vol}(g_{o})>v$ and ${\rm Ric}(g_{o})=(n-1)g_{o}$ satisfies $\lambda_{n-3}<n+\varepsilon_{0}$ with $\lambda_{k}$ the $k^{th}$ eigenvalue of the Laplace-Beltrami operator?
@@
Question 1.34 (Codimension-$5$ conjecture).
@@
Towards a proof in all dimensions. The present paper essentially proves the iteration from n1=4n-1=4 to $n=5$ of a plausible, long-term strategy in order to answer Question 1.34 affirmatively by an iteration on the dimension. Assume that noncollapsed limits of (n1)(n-1)-dimensional Einstein metrics have codimension $5$ regularity in the sense of Question 1.34.
@@ 1.4.2. Regularity of Einstein $5$-manifolds
Refined rectifiability. The isolation theorem implies that $1$-symmetry at a given scale at some point $x\in\mathcal{S}^{1}_{\Gamma}$ propagates to uniform symmetry across all smaller scales. Combined with the neck decomposition developed in [JN21, CJN21], this uniform symmetry allows one to identify the local $1$-stratum $S^{1}_{\Gamma}$ near $x$ with the center $\mathcal{C}_{0}$ of an associated neck region. A bi-Lipschitz coordinate is then constructed by using a splitting function to order points along $\mathcal{C}_{0}$ and pushing forward the Ahlfors regular measure associated to the neck region onto a subset of $\mathds{R}$ via the ordering. A countable selection of such parametrizations yields a covering of $\mathcal{S}^{1}$ by bi-Lipschitz images of subsets of $\mathbb{R}$, thereby establishing a refined rectifiability of the singular set without discarding any exceptional subsets. Moreover, we show that along each curve with constant volume density, the tangent cones remain identical. The proof is given in Section 4.3.
@@ 1.5. Acknowledgements
The first author is grateful to Tobias Colding for suggesting the uniqueness problem and for his patient guidance and many insightful discussions. The first author also thanks Wenshuai Jiang for his support and helpful conversations. The authors would like to thank Max Hallgren and Wenshuai Jiang for very valuable comments that helped us improve a step of the proof of Theorem 1.16 in Section 4.3.
During this project, Yiqi Huang was partially supported by a Simons Dissertation Fellowship and by NSF Grant DMS–2405393, and Tristan Ozuch was partially supported by the National Science Foundation under grant DMS–2405328.
@@ 4.1. Preliminaries
In [CJN21] and [JN21], they proved a structure theorem for the singular set $\mathcal{S}$:
@@
    If there exists a $\delta$-splitting function on $B_{2}(p)$, then $d_{GH}(B_{1}(p),B_{1}(0,x^{\prime}))\leqslant\varepsilon$ for $(0,x^{\prime})\in\mathds{R}\times X$. Moreover, $(u,v):B_{1}(p)\to\mathds{R}\times X$ gives an $\varepsilon$-Gromov-Hausdorff map for some $v:B_{1}(p)\to X$.
@@ 4.3. Refined rectifiability
Let $X^{5}$ be a noncollapsed Ricci-bounded limit space. Recall that in [CJN21, JN21], it is proven that the top stratum $\mathcal{S}^{1}\setminus\mathcal{S}^{0}$ is contained in a countable union of bi-Lipschitz images from subsets of $\mathds{R}$, modulo some $\mathcal{H}^{1}$-measure zero set. In this section, building on the uniqueness Theorem, we will improve the bi-Lipschitz structure by showing that there is no need to discard any measure zero subset, thereby establishing Theorem 1.16. For related work on uniqueness and the singular set we refer to [CM16, NV17, HJ23, HJ24, HJ25, FL25b, HKM25].
@@
First we prove a crucial uniform symmetry propagation result, coming from the gap Theorem 1.1.
@@
Next we will apply the uniform symmetry result to construct a local bi-Lipschitz map around each point in $\mathcal{S}^{1}_{\Gamma}$. Another key ingredient is the neck decomposition developed in [JN21, CJN21] (see also [NV19]).
We define the constant $\bar{V}=1/|\Gamma|$, the volume density at those points in $\mathcal{S}^{1}_{\Gamma}$. Given $x\in X$, $r>0$ and $\xi>0$, we define the volume-pinched set $\mathcal{P}_{r,\xi}(x)$ as
$\mathcal{P}_{r,\xi}(x)\equiv\{y\in B_{4r}(x):\mathcal{V}_{y}(\xi r)\leq\bar{V}+\xi\}.$
Most constants are chosen compatibly with the neck decomposition in [CJN21, Section 10], generally satisfying
$0<\xi<\delta<\gamma<\varepsilon<\tau\ll 1.$
@@
For any $p\in\mathcal{S}^{1}_{\Gamma}$, there exists some function $\phi$ defined on $B_{r}(p)$ for some radius $r>0$ such that $\phi$ is bi-Lipschitz on $B_{r}(x)\cap\mathcal{S}^{1}_{\Gamma}$: for some constant $C>1$ and any $x,y\in B_{r}(p)\cap\mathcal{S}^{1}_{\Gamma}$,
$C^{-1}\cdot d(x,y)\leqslant|\phi(x)-\phi(y)|\leqslant C\cdot d(x,y).$ (4.9)
@@
Fix $p\in\mathcal{S}^{1}_{\Gamma}$. After rescaling, by Proposition 4.13, there exists $R_{0}\gg 1$ and $\eta>0$ such that
$d_{GH}(B_{r}(x),B_{r}(0,x^{\prime}))\leq\eta r\text{ for any }x\in B_{4}(p)\cap\mathcal{S}^{1}_{\Gamma}\text{ and any }0<r\leq R_{0}$ (4.10)
with $(0,x^{\prime})$ the vertex of the cone $\mathds{R}\times\mathds{R}^{4}/\Gamma$. Moreover, we can assume $\mathcal{V}_{z}(\xi^{-1})\geq\bar{V}-\xi$ for any $z\in B_{4}(p)$ if $\eta$ is chosen sufficiently small.
We claim that if $d_{GH}(B_{r}(x),B_{r}(0,x^{\prime}))\leq\eta r$ with $\eta\leq\eta(v,\varepsilon,\gamma,\xi)$, then
${\rm Vol}\bigl(B_{\gamma r}(\mathcal{P}_{r,\xi}(x))\bigr)>\varepsilon\gamma^{4}r_{d}^{5}.$ (4.11)
Rescaling by $r_{d}$ and arguing by contradiction, we obtain a sequence of balls $B_{1}(x_{i})$ converging to $B_{1}(0,x^{\prime})\subset\mathds{R}\times\mathds{R}^{4}/\Gamma$ for which the above volume bound fails for fixed $v,\varepsilon,\gamma,\xi$. But $(-4,4)\times\{x^{\prime}\}\subset\mathcal{P}_{1,\xi}(0,x^{\prime})$ which satisfies ${\rm Vol}\bigl(B_{\gamma}(\mathcal{P}_{1,\xi}(0,x^{\prime}))\bigr)\geq c\,\gamma^{4}$ for some universal constant $c>0$. By volume convergence, this contradicts the failure of the bound whenever $\varepsilon<c/10$, proving the claim.
Hence we can apply the decomposition proposition for 1-symmetric balls by [CJN21, Proposition 10.5] (see also [JN21]), to obtain a decomposition of $B_{1}(p)$
$B_{1}(p)\subset\Big(\mathcal{C}_{0}\cup\mathcal{N}\cap B_{1}(x)\Big)\cup\bigcup_{b}B_{r_{b}}(x_{b})\cup\bigcup_{d}B_{r_{d}}(x_{d}),$
where
  1. (1)
    $\mathcal{N}=B_{2}(p)\setminus\Big(\mathcal{C}_{0}\cup\bigcup_{b}B_{r_{b}}(x_{b})\cup\bigcup_{d}B_{r_{d}}(x_{d})\Big)$ is a $(\delta,\tau)$-neck region;
  2. (2)
    each $B_{2r_{b}}(x_{b})$ satisfies $r_{h}(x_{b})>2r_{b}$, where $r_{b}$ is the harmonic radius;
  3. (3)
    each $B_{2r_{d}}(x_{d})$ satisfies ${\rm Vol}(B_{\gamma r_{d}}\mathcal{P}_{r_{d},\xi}(x_{d}))<\varepsilon^{2}\gamma^{4}r_{d}^{5}$ or $\mathcal{P}_{r_{d},\xi}(x_{d}))=\emptyset$.
The center set of the neck region is defined to be $\mathcal{C}=\mathcal{C}_{+}\cup\mathcal{C}_{0}$ where $\mathcal{C}_{+}=\bigcup_{b}\{x_{b}\}\cup\bigcup_{d}\{x_{d}\}$ with the corresponding radius. And if $x\in\mathcal{C}_{0}$, then $r_{x}=0$. Recall that the associated measure $\mu$ to the neck region $\mathcal{N}$ is defined by $\mu=\sum_{x\in\mathcal{C}_{+}}r_{x}\delta_{x}+\mathcal{H}^{1}|_{\mathcal{C}_{0}}$. By [CJN21, Theorem 2.9], for any $x\in\mathcal{C}$ and $r\geq r_{x}$ with $B_{2r}(x)\subset B_{2}(p)$, $\mu$ is Ahlfors regular: for some universal constant AA
$A^{-1}\cdot r\leq\mu(B_{r}(x))\leq A\cdot r.$ (4.12)
We claim that $\mathcal{C}_{0}\cap B_{2}(p)=\mathcal{S}^{1}_{\Gamma}\cap B_{2}(p)$. In fact, by cone-splitting 4.7 and $\epsilon$-regularity 4.8, the region $\mathcal{N}\cup\bigcup_{b}B_{r_{b}}(x_{b})$ is regular. It suffices to check that $B_{2r_{d}}(x_{d})\cap\mathcal{S}^{1}_{\Gamma}=\emptyset$. Suppose there exists some $x\in B_{2r_{d}}(x_{d})\cap\mathcal{S}^{1}_{\Gamma}$. Then by (4.10) and (4.11), we have $B_{r_{d}/2}(x)$ is $(1,\eta)$-symmetric. Therefore ${\rm Vol}(B_{\gamma r_{d}}\mathcal{P}_{r_{d},\xi}(x_{d}))>{\rm Vol}(B_{\gamma r_{d}/2}\mathcal{P}_{r_{d}/2,\xi}(x))>\varepsilon^{2}\gamma^{4}r_{d}^{5}$, contradicting the definition of the $d$-balls. This implies the claim.
Hence it suffices to find a function that is bi-Lipschitz on $\mathcal{C}_{0}$. By [CJN21, Proposition 9.3], there exists some $\delta^{\prime}$-splitting function $u$ defined on $B_{4}(p)$ such that for any $x,y\in\mathcal{C}$
$\frac{1}{2}\cdot d(x,y)^{2}\leq|u(x)-u(y)|\leq 2\cdot d(x,y).$
In particular, $u$ is 1-1 from $\mathcal{C}$ to some subset in $\mathds{R}$. We can now define the function
$\phi(x)\equiv\mu\Big(\{z\in B_{2}(p)\cap\mathcal{C}:u(z)\leq u(x)\}\Big).$
We will prove the bi-Lipschitz property for such $\phi$: for some universal constant $C>1$, and any $x,y\in\mathcal{C}_{0}$
$C^{-1}\cdot d(x,y)\leq|\phi(x)-\phi(y)|\leq C\cdot d(x,y).$
We fix two points $x,y\in\mathcal{C}_{0}$ with $d(x,y)=s$. Assume that $u(x)>u(y)$. Let $\Omega\equiv\{z\in B_{2}(p)\cap\mathcal{C}:u(y)<u(z)\leq u(x)\}$. Then $|\phi(x)-\phi(y)|=\mu(\Omega)$.
By [CJN21, Theorem 7.2], there exists some number $T=T_{x,4s}>0$ such that $Tu$ is a $\delta^{\prime\prime}$-splitting function on $B_{4s}(x)$. By Theorem 4.10, $B_{2s}(x)$ is $(1,\delta^{\prime\prime\prime})$-symmetric with respect to $\mathcal{L}_{x,2s}$ and $(Tu,v)$ is a $\delta^{\prime\prime\prime}$-GH map for some map $v:B_{2s}(p)\to\mathds{R}^{4}/\Gamma$. Hence it is easy to see that $\Omega\equiv\{z\in B_{4s}(x)\cap\mathcal{C}:Tu(y)<Tu(z)\leq Tu(x)\}$. This proves the upper bound for $\mu(\Omega)$ using (4.12) since in particular $\Omega\subset B_{4s}(x)$
$\mu(\Omega)\leq\mu(B_{4s}(x))\leq(4A)\cdot s.$ (4.13)
On the other hand, by the Reifenberg property of $\mathcal{C}$ (See [JN21, Definition 3.1 (n3)]), there exists some point $z\in B_{2s}(x)\cap\mathcal{C}$ such that $Tu|_{B_{s/10}(z)}\subset\big(Tu(y),Tu(x)\big)\subset\mathds{R}$. Moreover, by the Lipschitz property of the radius function on $\mathcal{C}$, we have $r_{z}\leq s/100$ since $r_{p}=r_{q}=0$. Therefore, we can apply (4.12) to obtain
$\mu(\Omega)\geq\mu(B_{s/10}(z))\geq(10A)^{-1}\cdot s.$ (4.14)
Combining (4.13) and (4.14) we complete the proof of the bi-Lipschitz property of $\phi$.
@@
Since we have the uniform lower bound of ${\rm Vol}(B_{1}(x))$ for $x\in B_{1}(p)$, the index set $I$ is finite. For each $x\in\mathcal{S}^{1}_{\Gamma_{i}}\cap B_{1}(p)$, by Proposition 4.14, there exists some radius $r$ such that $B_{r}(x)\cap\mathcal{S}^{1}_{\Gamma_{i}}$ is bi-Lipschitz to a subset of $\mathds{R}$. Since $X$ is separable, there exist countably many subsets $Z_{j}\in\mathcal{S}^{1}$ such that each $Z_{j}$ is bi-Lipschitz to some subset of $\mathds{R}$ with $\mathcal{S}^{1}=\bigcup_{j\in J}Z_{j}$. The proof is completed by noting that $\mathcal{S}^{0}$ is also countable. ∎